# Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting

### 03 Jan 2024 v1

@@ 3.3 Custom Kernels
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.Wen et al. (2023)
@@
, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values. Remark that in the context of self-attention, $Q=K=V=inputdata$.
@@
The Long Short-term memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$. Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$. The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$. The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state. The formulation of LSTM is as follows:
@@
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer    LogTrans
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
96 0.061 0.181 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208
192 0.085 0.22 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275
336 0.104 0.25 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302
720 0.149 0.303 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321
@@
Table 2: Univariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
The LSTM Kernel used in our experiment contains an LSTM cell and a linear layer that takes all hidden states of a sequence as input and gives a vector as output. The experimental result is shown in the section 4
4 Experiments and Results
Datasets.
We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Traffic and Electricity. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
Here, we followed the experiment setting in Zeng et al. (2023) and partitioned the data into 12, 4, and 4 months for training, validation, and testing respectively for the ETT dataset. The data is split into [0.7,0.1,0.2] for training, validation, and testing for dataset Weather, Traffic, and Electricity.
Baselines and Experimental Settings.
In the time series forecasting task, we follow the experiment setting in NLinear Zeng et al. (2023), which take $L={336,720}$ step historical data as input and forecast $T\in{96,192,336,720}$ step value in the future. We replace the last value normalization with mean value normalization for univariate forecasting in the ETT dataset. We include five recent methods: PatchTSTNie et al. (2023), NLinear, DLinearZeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021)and LogTrans Li et al. (2019). We merge the result reported in Nie et al. (2023) and Zeng et al. (2023) for taking their best in a supervised setting.
Following previous works Wu et al. (2021), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the core metrics to compare performance for Forecasting problems. We take a look back at Windows 336 and 720 in experiments. The list of multiples is respectively [3, 4, 4, 5] and [4, 4, 7] and the segment-unit input length is 3. The input dimension is 1 as we follow the channel-independent setting in Zeng et al. (2023). The hidden dimensions are 128 or 256 for multivariate tasks and 5 or 30 for univariate tasks depending on of dataset. We use the Random Erase Zhong et al. (2020) from image processing for data augmentation. The learning rate is selected in [0.00005, 0.001] depending on the dataset. The training epoch is 300 and the patience is 30 in general.
@@ 4 Experiments and Results
Instance Normalization and Mean Normalization Instance normalization Ulyanov et al. (2017) Kim et al. (2021) contains a step of substruction with the average value of the input sequence and a division of its standard deviation. It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model. Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises in the small dataset. We followed the mean normalization as the preprocessing step and compared it with instance normalization for selected models.
Model Variants.
We propose 4 kernels and 5 versions of kernel U-Net for experiments. The kernels tested are Linear, Linear with 1 hidden layer and non-linear activation, Transformer, and LSTM.
  • Linear U-Net is a baseline method that contains a list of linear kernels.
  • Linear-1-Hidden U-Net replaces 2 kernels close to latent vector to kernels with hidden layer and nonlinear activation.
  • Linear-5-Hidden U-Net replaces all linear kernels with non-linear kernels.
  • Transformer U-Net replaces 2 kernels close to latent vector to transformer kernels and keeps the rest as linear kernels.
  • LSTM U-Net replaces 2 kernels close to latent vector to LSTM kernels and keeps the rest as linear kernels.
The result shows that the performance of Kernel U-Net reaches or outperforms the state-of-art methods in univariate setting (Table 2) and reaches similar or better results in multivariate setting (Table 1).
5 Conclusion
In this paper, we propose Kernel U-Net, a symmetric and hierarchical framework that can effectively process long temporal dependencies with $O(log(L)^{2})$ number of parameters. Specifically, Kernel U-Net cut input sequences into small slices at each layer in the network for applying customized kernels.
In the future, we hope to develop more kernels that adapt well to particular datasets. We also hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.

### 12 Feb 2024 v2

@@ A.2 LSTM and Transformer kernels
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. We follow the description in Wen et al. [2023].
@@
Figure 8: Structure of Kernels a) Linear Kernel, b) Multi layer linear Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
@@
, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values.
@@
The transformer kernel in this work (Figure 8) follows the classical structure with 2 additional linear layers. The first linear layer before the Transformer block of size $M_{in}\times M_{out}$ reduces the dimension. The last linear layer is a matrix of size $L_{in}*M_{out}\times L_{out}*M_{out}$ that reduces the input length.
@@
The Long Short-term Memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$. Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$. The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$. The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state. The formulation of the LSTM cell is as follows:
@@
In our experiments, we add linear layers (Figure 8) that take the flattened output of all hidden states of the LSTM cell for linear combination.
$Z=W*Flatten([h_{1},...,h_{t}])+b$
, where $W\in R^{L_{i}*d_{h}}$, $b\in R^{L_{i+1}*M_{i+1}}$ and $d_{h}$ is the dimension of hidden state and $d_{h}=M_{i}$ in general. The residual skip operation aside from the LSTM cell contains a linear layer equivalent to a simple linear kernel discussed in the previous sections.
A.3 Dataset and Experiment
Dataset Statistics
We use 7 popular multivariate datasets for experiments. We create subset Weather(S), Electricity(S), and Traffic(S) from top-left to bottom-right for quick search of K-U-Net variants. The statistic is in noted in Table 2. We refer the description of datasets in Nie et al. [2023]:
  • Weather dataset collects 21 meteorological indicators in Germany, such as humidity and air temperature.
  • Traffic dataset records the road occupancy rates from different sensors on San Francisco freeways.
  • Electricity is a dataset that describes 321 customers’ hourly electricity consumption.
  • ETT (Electricity Transformer Temperature) datasets are collected from two different electric transformers labeled with 1 and 2, and each of them contains 2 different resolutions (15 minutes and 1 hour) denoted with m and h.
Table 2: Statistics of popular datasets for experiments.
Datasets ETTh1 ETTh2
ETTm1
ETTm2 Weather
Traffic
Electricity Weather(S) Traffic(S) Electricity(S)
Features 7 7
7
7 21
862
321 21 (100%) 21 (2.5%) 22 (7%)
Timesteps 17420 17420
69680
69680 52696
17544
26304 10539 (20%) 12280 (70%) 10521 (40%)
Table 3: Univariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best in underlined.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer LogTrans Yformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@ A.3 Dataset and Experiment
96 0.062 0.182 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208 0.128 0.275
192 0.089 0.224 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275 0.155 0.308
336 0.114 0.259 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302 0.208 0.358
720 0.156 0.312 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321 0.261 0.405
Table 4: Comparison of 16 variants of Linear K-U-Net in baseline setting and without using Skip connection(-Skip), Random Erase(-RE), Chanel Independence(-CI). The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined for minimum score(Min) and average score(Mean).
Method Min (Base) Min (-Skip) Min (-RE) Min (-CI) Mean (Base) Mean (-Skip) Mean (-RE) Mean (-CI)
Metric MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE
ETTh1
96 0.368 0.368 0.375 0.375 0.370 0.369 0.459 0.451 0.376 0.374 0.393 0.385 0.377 0.373 0.488 0.483
192 0.404 0.402 0.412 0.413 0.402 0.402 0.506 0.495 0.409 0.407 0.423 0.423 0.411 0.407 0.551 0.538
336 0.423 0.421 0.443 0.441 0.427 0.422 0.535 0.519 0.443 0.430 0.489 0.455 0.442 0.431 0.581 0.549
720 0.457 0.439 0.461 0.453 0.461 0.438 0.629 0.576 0.492 0.472 0.531 0.515 0.488 0.470 0.690 0.598
ETTh2
96 0.274 0.272 0.273 0.273 0.274 0.273 0.423 0.397 0.302 0.287 0.299 0.290 0.300 0.286 0.458 0.415
192 0.336 0.353 0.337 0.355 0.336 0.354 0.414 0.442 0.370 0.375 0.376 0.382 0.372 0.374 0.445 0.474
336 0.361 0.361 0.364 0.364 0.362 0.361 0.412 0.422 0.402 0.384 0.404 0.392 0.399 0.384 0.428 0.444
720 0.398 0.394 0.397 0.397 0.395 0.394 0.453 0.456 0.451 0.418 0.459 0.436 0.444 0.420 0.468 0.519
ETTm1
96 0.288 0.284 0.289 0.290 0.288 0.284 0.321 0.318 0.296 0.292 0.303 0.298 0.296 0.291 0.346 0.333
192 0.324 0.327 0.325 0.324 0.326 0.324 0.359 0.358 0.335 0.331 0.341 0.336 0.334 0.331 0.383 0.373
336 0.359 0.358 0.353 0.354 0.359 0.357 0.391 0.391 0.366 0.363 0.374 0.368 0.366 0.362 0.418 0.406
720 0.408 0.404 0.403 0.403 0.409 0.403 0.452 0.447 0.418 0.411 0.419 0.413 0.418 0.410 0.478 0.467
ETTm2
96 0.162 0.162 0.162 0.161 0.163 0.161 0.213 0.204 0.168 0.165 0.170 0.165 0.168 0.164 0.228 0.210
192 0.218 0.217 0.216 0.215 0.217 0.216 0.298 0.274 0.228 0.223 0.229 0.224 0.228 0.222 0.318 0.286
336 0.270 0.268 0.268 0.267 0.265 0.265 0.372 0.345 0.283 0.277 0.281 0.277 0.284 0.272 0.394 0.356
720 0.342 0.349 0.343 0.347 0.343 0.347 0.410 0.408 0.354 0.361 0.351 0.361 0.364 0.356 0.434 0.421
Average decrease +0.42% +0.77% +0.03% -0.20% +24.19% +21.45% +2.24% +2.30% +0.03% -0.32% +25.92% +23.72%
Experiment details
We use a 4-layer K-U-Net for experiments. The list of multiples are respectively [4, 4, 3, 7] and [4, 6, 6, 5] for look-back windows $L$ and forecasting horizons $T$ in {336, 720}. Remark that $T$ is fixed to be 336 for shorter forecasting cases. The bottom patch length is 4 and its width is 1 as we follow the Channel-Independent (CI) setting in Zeng et al. [2023]. In case of not using CI, the patch width equals to feature size. The hidden dimensions are 128 for multivariate and univariate tasks in general, except that they are 256 on the Traffic dataset. The learning rate is selected in $[0.00005,0.001]$ depending on the dataset. The training epoch is 50 and the patience of early stopping is 10 in general.
We apply weighted early stopping where the triggering MSE (TMSE) is computed with $TMSE_{e+1}=\tau*TMSE_{e}+(1-\tau)*VMSE_{e}$ , where $\tau$ is a hyper-parameter to be set, $e$ is the index of the episode, and VMSE is the running MSE on a validation set. The parameter $\tau$ sets the sensitivity of early stopping and is empirically selected. In general, it is set to be 0.9 in the variants searching stage and 0.5 in the fine-tuning stage.
The MSE loss function is defined with $:\mathcal{L}=\mathbb{E}_{\boldsymbol{x}}\frac{1}{B*M}\sum_{i=1}^{B*M}\left\|\hat{\boldsymbol{x}}_{L+1:L+T}^{(i)}-\boldsymbol{x}_{L+1:L+T}^{(i)}\right\|_{2}^{2}$ , where $B$ is the batch size and $M$ is the multiple of feature size unit of time series.
A.4 Search of model variants
We searched 30 variants of K-U-Net across Linear_[$0$,…,$1$] and {LSTM, Transformer}_[0100,…,$1$] for the ETT, Weather datasets and 3 subsets. The lookback window is typically 720, and we additionally tested a version of 336 on the ETT dataset. The result is reported with the average of running Top5 Minimum MSE (Top5MMSE) on the validation and test set in a single run. This measures the extreme performance of a variant and its potential in the fine-tuning stage. We use relative score (RS) to locate best-performed variants (i.g. $RS=1$ means that it’s the best among all candidates). We plot the RS of variants of 5 datasets in Figures 9 and Figures 10.
Based on the search result, We note that Linear and Multilayer linear kernels are more adapted in the ETT dataset than our Transformer and LSTM kernels. We remark that K-U-Net Linear_1000 is outstanding in the Electricity (S) dataset and thus selected for full examination.
A.5 Result of selected variants
In Table 5 6 7, we note selected models that are examined prudently with 5 runs in a fine-tuned context and measures with average MSE and MAE. We include the Table 9 showing that K-U-Net has a small standard deviation error across 5 random runs.
A.6 Result of univariate time series forecasting
The experiment of univariate time series forecasting task predicts the target feature is the ”oil temperature” on ETT datasets. We remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3). It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ in the Univariate setting compared to PatchTST.
Appendix B Ablation Study
For studying the roles of skip connection in K-U-Net, Random Erase, and Chanel Independence, we use the Minimum Top5 average MSE (Min Top5MMSE) to measure the extreme performance and the Mean Top5MMSE to measure the average performance. For instance Normalisation, we only show the comparison with Mean Normalization with the Traffic dataset because the improvement is evident.
@@ Appendix B Ablation Study
Table 5: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold.
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_0000
(720) K-U-Net
Linear_1100
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.372 0.393 0.377 0.401 0.370 0.407 0.370 0.393 0.370 0.400
192 0.405 0.414 0.413 0.424 0.404 0.433 0.404 0.414 0.413 0.429
336 0.431 0.430 0.444 0.444 0.420 0.445 0.420 0.430 0.422 0.440
720 0.438 0.454 0.457 0.471 0.482 0.490 0.438 0.454 0.447 0.468
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_0000
(720) K-U-Net
Linear_0001
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2
96 0.273 0.335 0.271 0.336 0.286 0.342 0.271 0.335 0.274 0.337
192 0.336 0.377 0.332 0.377 0.354 0.384 0.332 0.377 0.339 0.379
336 0.360 0.401 0.357 0.403 0.388 0.413 0.357 0.401 0.329 0.384
720 0.395 0.435 0.394 0.442 0.420 0.447 0.394 0.435 0.379 0.422
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_0000
(720) K-U-Net
Linear_1000
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.301 0.343 0.310 0.351 0.286 0.342 0.286 0.342 0.290 0.342
192 0.336 0.363 0.338 0.368 0.330 0.371 0.330 0.363 0.332 0.369
336 0.371 0.384 0.367 0.385 0.360 0.393 0.360 0.384 0.366 0.392
720 0.427 0.416 0.417 0.412 0.405 0.420 0.405 0.412 0.416 0.420
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_0000
(720) K-U-Net
Linear_0110
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm2
96 0.164 0.253 0.161 0.252 0.165 0.256 0.161 0.252 0.165 0.255
192 0.219 0.290 0.216 0.290 0.225 0.296 0.216 0.290 0.220 0.292
336 0.273 0.326 0.268 0.326 0.275 0.332 0.268 0.326 0.274 0.329
720 0.367 0.383 0.350 0.379 0.343 0.380 0.343 0.379 0.362 0.385
Table 6: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets Electricity, Weather and Traffic. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold.
Method K-U-Net
Linear_0000
(720) K-U-Net
Linear_1110
(720) K-U-Net
Linear_1000
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity
96 0.138 0.237 0.134 0.232 0.129 0.226 0.129 0.226 0.129 0.222
192 0.153 0.250 0.152 0.249 0.147 0.244 0.147 0.244 0.147 0.240
336 0.171 0.269 0.167 0.264 0.163 0.261 0.163 0.261 0.163 0.259
720 0.213 0.305 0.203 0.300 0.197 0.292 0.197 0.292 0.197 0.290
Method K-U-Net
Linear_0000
(720) K-U-Net
Transf_0001
(720) K-U-Net
LSTM_0001
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Traffic
96 0.412 0.280 0.399 0.262 0.358 0.253 0.358 0.253 0.360 0.249
192 0.425 0.287 0.415 0.278 0.373 0.262 0.373 0.262 0.379 0.250
336 0.438 0.294 0.424 0.295 0.390 0.271 0.390 0.271 0.392 0.264
720 0.458 0.314 0.465 0.309 0.430 0.292 0.430 0.292 0.432 0.286
Method K-U-Net
Linear_0110
(720) K-U-Net
Transf_0001
(720) K-U-Net
LSTM_0100
(720) MIN PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather
96 0.146 0.201 0.146 0.200 0.142 0.195 0.142 0.195 0.149 0.198
192 0.190 0.245 0.192 0.246 0.188 0.244 0.188 0.244 0.194 0.241
336 0.241 0.284 0.242 0.287 0.242 0.286 0.241 0.284 0.245 0.282
720 0.310 0.333 0.310 0.336 0.316 0.341 0.310 0.333 0.314 0.334
B.1 Instance Normalisation and Mean Normalisation
@@
Instance normalization contains a step of substruction with the average value of the input sequence and a division of its standard deviation. It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model. Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises into the dataset. We employed the mean normalization as the preprocessing step for most experiments except those with instance normalization on traffic datasets. The comparison in Table 8 shows that Instance Normalization improves the MSE by 3.89% for a K-U-Net LSTM_0001 on traffic datasets.
Table 7: Fine-tuned univariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold.
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_1111
(336) MIN PatchTST Nlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.053 0.177 0.054 0.182 0.053 0.177 0.055 0.179 0.053 0.177
192 0.069 0.205 0.079 0.219 0.069 0.205 0.071 0.215 0.069 0.204
336 0.076 0.222 0.076 0.221 0.076 0.221 0.076 0.22 0.081 0.226
720 0.077 0.224 0.078 0.226 0.077 0.224 0.087 0.232 0.08 0.226
Method K-U-Net
Linear_0110
(336) K-U-Net
Linear_1100
(336) MIN PatchTST Nlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2
96 0.114 0.267 0.121 0.272 0.114 0.267 0.129 0.282 0.129 0.278
192 0.146 0.306 0.161 0.317 0.146 0.306 0.168 0.328 0.169 0.324
336 0.157 0.321 0.178 0.340 0.157 0.321 0.171 0.336 0.194 0.355
720 0.192 0.355 0.182 0.347 0.182 0.347 0.223 0.38 0.225 0.381
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_0000
(720) MIN PatchTST Nlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.026 0.121 0.026 0.122 0.026 0.121 0.026 0.121 0.026 0.122
192 0.038 0.149 0.039 0.150 0.038 0.149 0.039 0.15 0.039 0.149
336 0.051 0.173 0.051 0.174 0.051 0.173 0.053 0.173 0.052 0.172
720 0.072 0.206 0.068 0.203 0.068 0.203 0.073 0.206 0.073 0.207
Method K-U-Net
Linear_0000
(336) K-U-Net
Linear_1000
(720) MIN PatchTST Nlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm2
96 0.063 0.182 0.062 0.183 0.062 0.182 0.065 0.187 0.063 0.182
192 0.090 0.224 0.089 0.227 0.089 0.224 0.093 0.231 0.09 0.223
336 0.116 0.259 0.114 0.263 0.114 0.259 0.121 0.266 0.117 0.259
720 0.170 0.319 0.156 0.312 0.156 0.312 0.171 0.322 0.17 0.318
Table 8: Comparison of Mean Normalization(+MN) and Instance Normalization (+IN) on Traffic dataset. We note the best results in bold..
Methods K-U-Net (+MN)
LSTM_0001
(720) K-U-Net (+IN)
LSTM_0001
Metric MSE MAE MSE MAE
Traffic
96 0.375 0.264 0.360 0.252
192 0.393 0.275 0.377 0.263
336 0.406 0.280 0.389 0.270
720 0.442 0.296 0.427 0.290
Table 9: MSE and standard deviation of K-U-Net across 5 random runs.
Method K-U-Net
LSTM_0001
(720) Method K-U-Net
LSTM_0100
(720) Method K-U-Net
Linear_1000
Metric MSE Metric MSE Metric MSE
Traffic
96 0.35761 ±0.00042
Weather
96 0.14200 ±0.00016
ETTm1
96 0.28600 ±0.00056
192 0.37266 ±0.00056 192 0.18782 ±0.00078 192 0.32996 ±0.00204
336 0.39030 ±0.00135 336 0.24164 ±0.00139 336 0.35979 ±0.00092
720 0.42970 ±0.00078 720 0.31646 ±0.00198 720 0.40539 ±0.00128
B.2 Skip Connection,
Skip Connection in the U-Net is used for passing information from the encoder to the same level on its symmetric decoder. As the size of the hidden vector is reduced during passage in the encoder, skip connection helps keep the low-level features and allows the decoder to mix with high-level information. We conduct experiments to understand the role of skip connection in the K-U-Net. The result in Table 4 shows that removing the skip connection degrades the Min Top5MMSE score for $0.77\%$ on average and especially $3.19\%$ on the ETTh1 dataset. Additionally, it degrades $2.30\%$ Mean Top5MMSE on average and especially $5.42\%$ on the ETTh1 dataset.
B.3 Comparison of Chanel Independent and Chanel Mixing
Chanel Independent was an inherited choice from Zeng et al. [2023]. It avoids introducing noises that degrade the performance across different features. In this experiment, we compare the channel-independent(CI) strategy and the channel-mixing strategy. In K-U-Net, setting input and output dimension to 1 implies a channel-independent strategy, while setting to feather size (7 for the ETT dataset) means a channel-mixing strategy. As We observe in Table 4, removing CI degrades the Min Top MMSE score for $21.45\%$ and the Min Top5MMSE score for $23.72\%$.
B.4 Comparison of Random Erasing
Random Erasing augments the dataset by masking partially the input. In this experiment, we studied the role of random erasing in time series forecasting settings. We observe in Table 4 that removing Random Erasing degrades the Min Top5MMSE score for $0.2\%~{}0.32\%$, which is not as evident as the Skip Connection and Channel Independent strategy.
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Weather
(f) Weather(S)
(g) Electricity(S)
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Weather
(f) Weather(S)
(g) Electricity(S)
(h) Traffic(S)
Figure 9: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on validation and test set.
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Weather
(f) Weather(S)
(g) Electricity(S)
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Weather
(f) Weather(S)
(g) Electricity(S)
(h) Traffic(S)
Figure 10: Search result of K-U-Net with Multi-layer Linear kernels on validation and test set.